OpenStudy (aaronandyson):

3x= 4h 5x + 960 = 12h Find h and x

OpenStudy (babynini):

Hello :) do you have an idea of how to work this?

OpenStudy (babynini):

ooo those are not pleasant numbers D:

OpenStudy (aaronandyson):

? 5x + 960 = 12x x = 240

OpenStudy (aaronandyson):

@Sepeario

OpenStudy (babynini):

Aaron you got your equation wrong.

OpenStudy (babynini):

:P

OpenStudy (aaronandyson):

?

OpenStudy (babynini):

5x + 960 = 12x x = 240 but it's 5x +960 = 12h

OpenStudy (aaronandyson):

3x = 4h Can I multiply by 3?

OpenStudy (babynini):

um yeah in that case you could but those aren't the numbers

OpenStudy (aaronandyson):

@perl

OpenStudy (babynini):

haha yeah idk how to help with this one D: sorry!

OpenStudy (aaronandyson):

:(

OpenStudy (babynini):

@rvc

OpenStudy (aaronandyson):

240 = x is correct :P

OpenStudy (aaronandyson):

They said it x = 240 h = 180 XD